# Imbalanced Data Clustering using Equilibrium K-Means

### v1

@@ V Numerical Experiments
Numerical experiments are conducted to compare the performance of our proposed EKM algorithm with six other algorithms: (1) HKM, (2) FKM, (3) MEFC, (4) csiFKM [19], (5) siibFKM [20], and (6) the Gaussian mixture model. Multiprototype clustering algorithms are not used as benchmarks, as the multiprototype mechanisms can be combined with EKM for fair comparison. This is beyond the scope of this paper because the paper does not intend to find the best KM algorithm to pair with the multiprototype mechanism.
To avoid local optima convergence, each clustering algorithm is executed ten times, presenting the lowest objective value result. Centroids are initialized independently for each replication using the K-means++ algorithm [33]. Convergence is deemed achieved when the relative moving distance of centroids between successive iterations is less than $1\mathrm{e}{-3}$, i.e.,
@@
The maximum number of iterations is set to 100. FKM, csiFKM, and siibFKM employ a typical fuzzifier value of $m=2$ [34]. The fuzzifier value of MEFC is set to $\lambda=1$ and the smoothness value of EKM is set to $\alpha=1$. We test the algorithms on 12 datasets, including two synthetic and ten real datasets. Because of space limitations, only partial experiment results from six datasets are shown in the paper. Full experiment results with quantitative metrics can be found in the supplemental material. The six datasets are introduced below.
@@ V-A1 Synthetic 2-D Ball
We create a synthetic dataset with data independently generated from three distinct normal distributions, each having different means: $(-2,2)$, $(2,-2)$, and $(4,4)$. Their covariance matrices are all identity matrices. The dataset consists of 2,000 samples from the normal distribution with means $(-2,2)$, and 50 samples from the other two normal distributions, resulting in a total of 2,100 samples. The clustering results are shown in Fig. 4.
@@ V-A2 Synthetic 2-D Noisy Ball
We add 100 uniformly distributed noise data samples (around 5% of the number of clean samples) to the synthetic 2-D ball dataset to test the algorithms’ robustness against noise. The clustering outcomes are displayed in Fig. 5.
@@ V-A4 Imbalanced Fisher’s Iris Data
Because the instance distribution in the Iris dataset is balanced. To examine the algorithms’ capability for imbalanced clustering, we generate an imbalanced Fisher’s Iris data by removing the first 30 instances of Iris setosa, and merging the other two species (Iris virginica and Iris versicolor) into a single class. The clustering results for this dataset are presented in Fig. 7.
@@ V-A5 Wisconsin Diagnostic Breast Cancer
Wisconsin Diagnostic Breast Cancer (WDBC) [37] comprises 30 features of cell nuclei from diagnostic benign and malignant breast tumors, obtained from digitized images of fine-needle aspiration of breast masses. This database contains 569 instances, with a fairly even distribution of 357 benign and 212 malignant cases. We utilize the first three features (mean radius, mean texture, and mean perimeter) for clustering, each normalized with zero mean and unit variance. Fig. 9 presents the clustering results where the mean radius and the mean texture are used as coordinates for visualization.
@@ V-A6 Imbalanced Wisconsin Diagnostic Breast Cancer
The data distribution in WDBC is quite balanced. To evaluate the imbalanced data clustering performance, we remove the first 200 malignant instances in WDBC, resulting in an imbalanced WDBC database with a class distribution of 357 benign and 12 malignant instances. We still use the first three features for clustering, each normalized with zero mean and unit variance. The clustering results are shown in Fig. 9.
V-A7 ColoredChips
This is not a dataset but an image named coloredChips.png from Matlab’s image gallery (shown in Fig. 10a). The purpose of this experiment is to test the algorithms’ performance on the task of color-based image segmentation. The number of clusters is set to five to expect the segmentation of the four colored chips and the background (i.e., the desk surface). Clustering is conducted in the RGB space and the segmentation results can be viewed in Fig. 10b - 10e.
@@ V-B Discussion of Results
In general, these results demonstrate that the proposed EKM algorithm performs competitively on balanced data and outperforms the other clustering algorithms on imbalanced data. In particular, HKM, FKM, and MEFC cannot find the correct clusters in imbalanced datasets. The csiFKM and siibFKM algorithms are sensitive to noise data. Fig. 5c and Fig. 5d show that their centroids are highly deviated from the correct positions because of noise. In comparison, EKM is more robust, as seen in Fig. 5e.
Fig. 9 displays the clustering results of HKM, FKM, GMM, and EKM for the imbalanced WDBC dataset. The strong performance of EKM indicates its potential application in anomaly detection, particularly in identifying clustered anomalies. Although HKM and FKM are also used for anomaly detection, they typically detect isolated anomalies, not clustered ones [38, 39, 40, 41].
The color-based image segmentation results are shown in Fig. 10b - 10e. It can be seen that HKM, FKM, and MEFC divide the background (the desk surface) into two clusters. This happens because the background comprises most pixels, and splitting it into two clusters better balances the number of pixels within the clusters. On the other hand, EKM produces the expected segmentation result.
@@
Figure 10: (a) coloredChips.png (b) Segmentation by HKM. (c) Segmentation by FKM. (d) Segmentation by MEFC. (e) Segmentation by EKM.
@@ VI Deep Clustering
Clustering algorithms using Euclidean distance for defining similarity often face challenges when clustering high-dimensional data, such as images, due to the small distance differences between various point pairs in high-dimensional space [42]. Deep clustering is a technique that addresses this issue by employing DNNs to map high-dimensional data non-linearly to low-dimensional representation, making the subsequent clustering easier. Most existing deep clustering approaches combine DNNs with HKM, e.g., [12, 13, 14]. However, this design is less favorable to imbalanced datasets [13]. In this section, we provide an example of replacing HKM with EKM in a popular deep clustering framework, showing its performance improvement on an imbalanced dataset derived from the MNIST.
@@ VI-A Optimization Procedure
Deep clustering network (DCN) [12] is a popular deep clustering framework, as illustrated in Fig. 11. Essentially, DCN maps high-dimensional data to a low-dimensional representation through an autoencoder network. An autodecoder follows the autoencoder, mapping the representation back to the original high-dimension space (i.e., reconstruction). To ensure that the low-dimensional representation maintains the primary information of the original data, the autoencoder and autodecoder are jointly trained to minimize the reconstruction error. Additionally, to create a clustering-friendly structure in the low-dimensional representation, a clustering error is minimized along with the reconstruction error. DCN uses an alternating optimization algorithm to minimize the total error, and the optimization process is described below.
@@
where $m_{k}^{n}$ is the number of samples assigned to the $k$-th cluster before the incoming data $\mathbf{x}_{n}$, controlling the learning rate of the $k$-th centroid. Overall, the optimization procedure of DCN with HKM alternates between updating networks parameters $\{\bm{\theta}_{e},\bm{\theta}_{d}\}$ by solving (45) and updating HKM parameters $\{\mathbf{C},\{\mathbf{s}_{n}\}\}$ by (46) and (47).
@@
where $d_{kn}=\frac{1}{2}\|\bm{f}(\mathbf{x}_{n})-\mathbf{c}_{k}\|_{2}^{2}$. There are other details and tricks to implement DCN, such as the initialization of the networks. We only introduce the part related to our contribution, and for the rest, the readers can refer to the original paper of DCN [12].
@@ VI-B Clustering Performance on MNIST
To implement DCN, we refer to the code provided by one of its authors, available at https://github.com/boyangumn/DCN-New. We use the default neural network structure and hyperparameters. In particular, the deep representation dimension is set to ten, and the parameter $\beta$ is set to one. The smoothness parameter $\alpha$ for EKM is determined by the strategy introduced in Section IV-C. We evaluate clustering performance using three widely accepted metrics: including normalized mutual information (NMI), adjusted Rand index (ARI), and clustering accuracy (ACC). NMI and ACC range from 0 to 1, with zero representing the worst and one representing the best. ARI ranges from -1 to 1, with minus one representing the worst and one representing the best.
We first validate the algorithms on the full MNIST dataset [43], containing 70,000 gray images of handwritten digits from 0 to 9. We set the smoothness parameter of EKM to $\alpha=5\mathrm{e}-3$. The clustering results are presented in TABLE I, where we compare the proposed DCN+EKM with DCN+HKM and stacked autoencoder (SAE). SAE is a specific version of DCN that only minimizes the reconstruction error. We omit the results of DCN+FKM and DCN+MEFC due to space limitations and their poorer performance compared to DCN+HKM. The results show that DCN outperforms SAE, highlighting the importance of a clustering-friendly representation. The table also reveals that DCN+EKM performs similarly to DCN+HKM for this balanced dataset.
To evaluate the performance of the algorithms for imbalanced data clustering, we remove the MNIST dataset’s training set samples with numbers 1 to 9 (leaving only the test set samples). This imbalanced dataset has 15,923 images, with around 6,900 being the number 0 and the remaining numbers (1 to 9) being around 1,000. We set the smoothness parameter of EKM to $\alpha=3.8\mathrm{e}-3$. The clustering results are summarised in TABLE II, showing that DCN+EKM significantly outperforms DCN+HKM. To analyze the reason, we use t-SNE [44] to map the ten-dimensional representation obtained by DCN to a two-dimensional space for visualization, with the results displayed in Fig. 12. We observe that the deep representation obtained by DCN+EKM is more discriminative than DCN+HKM, and DCN+EKM successfully separates the large class (number 0) from other small classes, while DCN+HKM erroneously groups the large class into four clusters.
@@
Figure 12: Visualization of low-dimensional DCN representations by t-SNE. (a) Representations obtained by DCN+HKM, colored by true labels. (b) Representations obtained by DCN+HKM, colored by clusters of HKM. (c) Representations obtained by DCN+EKM, colored by true labels. (d) Representations obtained by DCN+EKM, colored by clusters of EKM.
@@ VII Conclusion
This paper presents equilibrium K-means (EKM), a novel clustering algorithm that is effective for imbalanced data. EKM is simple, interpretable, scalable to large datasets, and capable of finding unequal-sized clusters, making it a better alternative for hard K-means (HKM) and fuzzy K-means (FKM) for imbalanced data. Experimental results on real datasets from various domains demonstrate that EKM outperforms HKM and FKM on datasets with imbalanced data and performs comparably on datasets with balanced data. Furthermore, we provide a general framework that unifies HKM, FKM, and EKM. We encourage readers to explore and develop new K-means algorithms based on this more flexible framework to address other specific K-clustering problems besides imbalanced data distribution.
@@ Appendix A Proof of Theorem 1
where $\epsilon$ is a positive number. Hence, the sequence $(J^{(1)},J^{(2)},\cdots)$ is non-increasing, and with the boundness condition that $h>-\infty$, we have $\lim_{\tau\to+\infty}(J^{(\tau)}-J^{(\tau+1)})\to 0$. If the left side of the inequality (51) converges to zero, the right side of the inequality also converges to zero since it is non-negative. Consequently, we have $\lim_{\tau\to+\infty}\partial_{\mathbf{c}_{k}}J^{(\tau)}\to 0$ for all $k$. Therefore, the sequence $(\mathbf{c}_{k}^{(\tau)})$ converges to a stationary point of the objective function $J$. Because $(J^{(1)},J^{(2)},\cdots)$ is non-increasing, only (local) minimizers or saddle points appear as limit points.
@@
We evaluate the clustering performance of the proposed Equilibrium K-means (EKM) on ten real datasets. Four datasets (Iris, Imbalanced Iris, Wisconsin Diagnostic Breast Cancer (WDBC), and Imbalanced WDBC) are discussed in detail in the main text. This supplemental material presents the results for the remaining six datasets from the UCI repository (available at https://archive.ics.uci.edu/datasets): Wine, Ecoli, Image Segmentation (IS), Htru2, Rice, and Dry Bean. A detailed description of all ten datasets can be found in TABLE SI, ordered according to the imbalance level measured by CV0, which we will define later. Datasets with a lower order have higher CV0 values, indicating a higher imbalance level. We provide some necessary explanations for this table below.
In the “Feature” column, the format “x (y)” indicates that the dataset has y features, with x features used for clustering. For instance, the Iris dataset has four features, with two used for clustering. Only the four datasets in the main text use partial features for clustering, which serves the purpose of visualizing clustering results. For the additional six datasets, all features are used since visualization is not required. In the “CV0” column, we calculate the coefficient of variation (CV) to measure data imbalance. CV is defined as the ratio of the standard deviation to the mean. CV is defined as the ratio of the standard deviation to the mean. Mathematically, given the number of instances in each class as $N_{1},\cdots,N_{K}$, we have
@@
Each dataset is normalized so that each feature has zero mean and unit variance. The number of clusters is assumed to be equal to the number of classes for all datasets. For the supplementary six datasets, the smoothness parameter $\alpha$ is chosen as follows:
@@
where $\bar{d}_{0n}=\frac{1}{2}\sum_{n=1}^{N}\|\mathbf{x}_{n}\|_{2}^{2}/N$, $\mathbf{x}_{n}$ is the $n$-th data point, and $N$ is the total number of data points. The value $\bar{d}_{0n}$ can be interpreted as a data variance after the normalization of zero data mean. The normalized mutual information (NMI), adjusted Rand index (ARI), and clustering accuracy (ACC) metrics of each algorithm on the ten datasets are presented in TABLE SII. Bold values indicate the best results. We also include three more metrics: “CV1”, ”DCV” and ”Running time” for readers’ interest. “CV1” is the CV value of the clustering result. “DCV” is the difference between CV0 and CV1, i.e., $\text{DCV}=\text{CV}0-\text{CV}1$. A positive DCV value indicates that the data distribution in the true classes is more imbalanced than in the clusters, while a negative DCV value means the data distribution in the true classes is more uniform than in the clusters. “Running time” is the sum of the running time of ten replicates.
We observe that the proposed EKM generally outperforms the other six algorithms on the last four datasets with CV values greater than 0.5. EKM outperforms other algorithms on two of the top six datasets with CV values less than 0.5. On the rest of the datasets, EKM still has a competitive performance. The running time of EKM is also comparable to FKM. Overall, these results demonstrate that EKM is an efficient algorithm for clustering balanced and imbalanced data.

### v2

@@ V Numerical Experiments
Numerical experiments are conducted to compare the performance of our proposed EKM algorithm with six other centroid-based algorithms: (1) HKM, (2) FKM, (3) MEFC, (4) csiFKM [19], (5) siibFKM [20], and (6) the GMM. Multiprototype-based algorithms such as [22, 5] are not appropriate as baseline models because they are too complex to be benchmarks for gauging the efficiency of the proposed EKM.
To avoid local optima convergence, each clustering algorithm is executed ten times, and the result with the lowest objective value is presented. Centroids are initialized independently for each replication by the K-means++ algorithm [32]. Convergence is achieved when the relative moving distance of centroids between successive iterations is less than $1\mathrm{e}{-3}$, i.e.,
@@
The maximum number of iterations is set to 100. FKM, csiFKM, and siibFKM employ a typical fuzzifier value of $m=2$ [33]. The fuzzifier value of MEFC is set to $\lambda=1$ and the smoothing parameter of EKM is set to $\alpha=1$. We evaluate EKM’s performance on 12 datasets, including two synthetic and ten real datasets. Besides, we also apply EKM to the application of image segmentation. Because of space limitations, experiment results from six datasets and the image segmentation are shown in the paper. Full experiment results can be found in the supplemental material. The six datasets and the test image are introduced below.
@@ V-A1 Synthetic 2-D Ball
We generate data from three distinct normal distributions, each having different means: $(-2,2)$, $(2,-2)$, and $(4,4)$, and the same identity covariance matrix. The dataset consists of 2,100 samples, of which 2000 samples from the normal distribution with means $(-2,2)$, and 100 samples from the other two normal distributions (each distribution for 50 samples). The clustering results are shown in Fig. 4.
@@ V-A2 Synthetic 2-D Noisy Ball
We add 100 uniformly distributed noise data (around 5% of the number of clean samples) to the synthetic 2-D ball dataset in order to test the algorithms’ robustness against noise. The clustering outcomes are displayed in Fig. 5.
@@ V-A4 Imbalanced Fisher’s Iris Data
Because the instance distribution in the Iris dataset is balanced. To examine the capability of imbalanced data clustering, we generate an imbalanced Fisher’s Iris dataset by removing the first 30 instances of Iris setosa, and merging the other two species (Iris virginica and Iris versicolor) into a single class. The clustering results of this imbalanced dataset are presented in Fig. 7.
@@ V-A5 Wisconsin Diagnostic Breast Cancer
Wisconsin Diagnostic Breast Cancer (WDBC) [36] comprises 30 features of cell nuclei from diagnostic benign and malignant breast tumors, obtained from digitized images of fine-needle aspiration of breast masses. This database contains 569 instances, with an instance distribution of 357 benign and 212 malignant cases. We use the first three features (mean radius, mean texture, and mean perimeter) for clustering, each normalized with zero mean and unit variance. Fig. 8 presents the clustering results where the mean radius and the mean texture are used as coordinates for visualization.
@@ V-A6 Imbalanced Wisconsin Diagnostic Breast Cancer
The data distribution in WDBC is quite balanced. To examine the ability of imbalanced data clustering, we remove the first 200 malignant instances in WDBC, resulting in an imbalanced WDBC database with a distribution of 357 benign and 12 malignant instances. We still use the first three features for clustering, each normalized with zero mean and unit variance. The clustering results are shown in Fig. 9.
V-A7 Color-based image segmentation
The test image coloredChips.png is chosen from Matlab’s image gallery (shown in Fig. 10a). The number of clusters is set to five to expect the segmentation of the four colored chips and the background (i.e., the desk surface). Clustering is conducted in the RGB color space and the segmentation results can be viewed in Fig. 10b - 10e.
@@ V-B Discussion of Results
In general, these results demonstrate that the proposed EKM algorithm performs competitively on balanced data and outperforms the other clustering algorithms on imbalanced data. In particular, HKM, FKM, and MEFC suffer from the uniform effect when data has an imbalanced distribution. The csiFKM and siibFKM algorithms are sensitive to noise data: Fig. 5c and Fig. 5d show that their centroids are highly deviated from the correct positions in the presence of noise. In comparison, EKM is more robust to noise because of the anchoring effect produced by the positive gradients (see Section IV-B), as seen in Fig. 5e.
The strong performance of EKM shown in Fig. 9 demonstrates its ability to identify clustered anomalies. This makes up a missing function of HKM and FKM in anomaly detection: they are typically used to detect isolated anomalies, not clustered anomalies [37, 38, 39, 40].
The results of color-based image segmentation are shown in Fig. 10b - 10e. HKM, FKM, and MEFC erroneously separate the background (the desk surface) into two clusters. This happens because the background comprises most pixels, and splitting it into two clusters better balances the number of pixels within each clusters. On the other hand, EKM produces the expected segmentation result.
@@
Figure 10: (a) coloredChips.png. (b) Segmentation by HKM. (c) Segmentation by FKM. (d) Segmentation by MEFC. (e) Segmentation by EKM.
@@ VI Deep Clustering
Clustering algorithms using Euclidean distance for defining similarity face a challenge when clustering high-dimensional data, such as images, due to the small distance differences between various point pairs in high-dimensional space [41]. Deep clustering is a technique that addresses this issue by employing DNNs to map high-dimensional data to low-dimensional representation. Many deep clustering methods combine DNNs with HKM, e.g., [15, 16, 17]. However, this design is less favorable to imbalanced data [16]. In this section, we show the promise of EKM in the deep clustering of imbalanced data by replacing HKM with EKM in a popular deep clustering framework.
@@ VI-A Optimization Procedure
Deep clustering network (DCN) [15] is a popular deep clustering framework. As illustrated in Fig. 11, DCN maps high-dimensional data to low-dimensional representation through an autoencoder network. An autodecoder follows the autoencoder, mapping the representation back to the original high-dimensional space (i.e., reconstruction). To ensure that the low-dimensional representation maintains the primary information of the original data, the autoencoder and the autodecoder are jointly trained to minimize the reconstruction error. Additionally, to make the low-dimensional representation have a clustering-friendly structure, a clustering error is minimized along with the reconstruction error. DCN uses an alternating optimization algorithm to minimize the total error, and the optimization process is described below.
@@
where $m_{k}^{n}$ is the number of samples assigned to the $k$-th cluster before the incoming data $\mathbf{x}_{n}$, controlling the learning rate of the $k$-th centroid. Overall, the optimization procedure of DCN alternates between updating networks parameters $\{\bm{\theta}_{e},\bm{\theta}_{d}\}$ by solving (45) and updating HKM parameters $\{\mathbf{C},\{\mathbf{s}_{n}\}\}$ by (46) and (47).
@@
where $d_{kn}=\frac{1}{2}\|\bm{f}(\mathbf{x}_{n})-\mathbf{c}_{k}\|_{2}^{2}$. There are other details and tricks to implement DCN, such as the initialization of the networks. We only introduce the part related to our contribution, and kindly refer to [15] for more implementation details.
@@ VI-B Clustering Performance on MNIST
To implement DCN, we refer to the code one of its authors provided, available at https://github.com/boyangumn/DCN-New. We use the default neural network structure and hyperparameters. In particular, the dimension of the low-dimensional representation is set to ten, and the parameter $\beta$ is set to one. The smoothing parameter $\alpha$ is tuned on the representation obtained by the initialized DCN networks according to the strategy introduced in Section IV-C. We evaluate clustering performance using three widely accepted metrics: including normalized mutual information (NMI), adjusted Rand index (ARI), and clustering accuracy (ACC). NMI and ACC range from 0 to 1, with zero representing the worst and one representing the best. ARI ranges from -1 to 1, with minus one representing the worst and one representing the best.
We first evaluate the algorithm’s performance on the full MNIST dataset [42], which contains 70,000 gray images of handwritten digits from 0 to 9. Each image has $28\times 28=784$ pixels. We set the smoothing parameter of EKM to $\alpha=5\mathrm{e}-3$. The clustering results are presented in TABLE I. We compare the proposed DCN+EKM with DCN+HKM and stacked autoencoder (SAE). SAE is a specific version of DCN that only minimizes the reconstruction error. Thus, the learned representation by SAE does not have a clustering-friendly structure. The results show that DCN outperforms SAE, highlighting the importance of a clustering-friendly structure. We can also see that DCN+EKM performs similarly to DCN+HKM. We omit the results of DCN+FKM and DCN+MEFC due to space limitations. Their performance is not better than DCN+HKM.
MNIST contains 60,000 training images and 10,000 testing images. To evaluate the performance of imbalanced data clustering, we generate an imbalanced MNIST dataset by removing the training images of digits 1 to 9. This imbalanced dataset contains 15,923 images, of which approximately 6,900 are digit 0, while the remaining digits (1 to 9) each have about 1,000 images. We set the smoothing parameter of EKM to $\alpha=3.8\mathrm{e}-3$. The results are summarized in TABLE II. In conclusion, we find that an EKM-friendly representation structure is crucial for deep clustering of imbalanced data. We map the ten-dimensional representation obtained by DCN to a two-dimensional space by t-SNE [43] for visualization. The results are displayed in Fig. 12. The visualization result implies that the deep representation obtained by DCN+EKM is more discriminative than that obtained by DCN+HKM. We observe that EKM successfully identifies the large class (number 0) from other small classes without referring to the true labels, while HKM incorrectly divides the large class into four clusters.
@@
Figure 12: Visualization of DCN representations by t-SNE. (a) Representations obtained by DCN+HKM, colored by true labels. (b) Representations obtained by DCN+HKM, colored by labels generated by HKM. (c) Representations obtained by DCN+EKM, colored by true labels. (d) Representations obtained by DCN+EKM, colored by labels generated by EKM.
@@ VII Conclusion
This paper presents equilibrium K-means (EKM), a novel clustering algorithm effective for both balanced and imbalanced data. EKM is simple, interpretable, scalable to large datasets, and capable of creating unequal-sized clusters. Experimental results on real datasets from various domains show that EKM outperforms HKM, FKM, and other popular centroid-based algorithms on datasets with imbalanced data and performs comparably on datasets with balanced data. In the same deep clustering framework, compared with HKM and FKM, the clustering accuracy of imbalanced data using EKM has been noticeably improved. Furthermore, we reformulate HKM, FKM, and EKM in a general form of gradient descent. We encourage readers to study more properties of this general form to develop new K-means algorithms with non-heuristic memberships to address different issues.
@@ Appendix A Proof of Theorem 1
where $\epsilon$ is a positive number. Hence, the sequence $(J^{(1)},J^{(2)},\cdots)$ is non-increasing, and with the boundness condition that $h>-\infty$, we have $\lim_{\tau\to+\infty}(J^{(\tau)}-J^{(\tau+1)})\to 0$. If the left side of the inequality (LABEL:theorem1_proof_inequality) converges to zero, the right side of the inequality also converges to zero since it is non-negative. Consequently, we have $\lim_{\tau\to+\infty}\partial_{\mathbf{c}_{k}}J^{(\tau)}\to 0$ for all $k$. Therefore, the sequence $(\mathbf{c}_{k}^{(\tau)})$ converges to a stationary point of the objective function $J$. Because $(J^{(1)},J^{(2)},\cdots)$ is non-increasing, only (local) minimizers or saddle points appear as limit points.
@@
We evaluate the clustering performance of the proposed equilibrium K-means (EKM) on ten real datasets. Four datasets (Iris, Imbalanced Iris, Wisconsin Diagnostic Breast Cancer (WDBC), and Imbalanced WDBC) are introduced in detail in the main text. Herein we supplement the results for the remaining six datasets. They are from the UCI repository (available at https://archive.ics.uci.edu/datasets), which are: Wine, Ecoli, Image Segmentation (IS), Htru2, Rice, and Dry Bean. A detailed description of all ten datasets are summarized in TABLE SI, ordered according to the imbalance level measured by CV0, which we will define later. The dataset with a lower order has a higher CV0 value, indicating a higher imbalance level. In the “Feature” column of TABLE SI, the format “x (y)” indicates that the dataset has y features, of which x features are used for clustering. For instance, the Iris dataset has four features, of which two are used for clustering. Partial features are used for the sake of visualizing clustering results. In the “CV0” column, we calculate the coefficient of variation (CV) of the dataset to measure the level of data imbalance. CV is defined as the ratio of the standard deviation to the mean. Mathematically, given the number of instances in each class as $N_{1},\cdots,N_{K}$, we have
@@
Each dataset is normalized so that each feature has zero mean and unit variance. The number of clusters is assumed to be equal to the number of classes for all datasets. For the supplementary six datasets, the smoothness parameter $\alpha$ is chosen adaptively as follows:
@@
where $\bar{d}_{0n}=\frac{1}{2}\sum_{n=1}^{N}\|\mathbf{x}_{n}\|_{2}^{2}/N$, $\mathbf{x}_{n}$ is the $n$-th data point, and $N$ is the total number of data points. The value $\bar{d}_{0n}$ can be interpreted as a data variance after the normalization of zero data mean. The normalized mutual information (NMI), adjusted Rand index (ARI), and clustering accuracy (ACC) of each algorithm on the ten datasets are presented in TABLE SII. Bold values indicate the best results. We also include three metrics: “CV1”, ”DCV” and ”Running time” for readers’ interest. “CV1” is the CV value of the clustering result. “DCV” is the difference between CV0 and CV1, i.e., $\text{DCV}=\text{CV}0-\text{CV}1$. A positive DCV value indicates that the data distribution of the true classes is more imbalanced than that of the generated clusters, while a negative DCV value means the data distribution of the true classes is more uniform than that of the generated clusters. “Running time” is the sum of the running time of ten replicates.
We observe that the proposed EKM generally outperforms the other six algorithms on the last four datasets with CV values greater than 0.5. EKM outperforms other algorithms on two of the top six datasets with CV values less than 0.5. On the rest of the datasets, EKM still has a competitive performance. Overall, these results demonstrate that EKM is an efficient clustering algorithm for both balanced and imbalanced data.